OpenStudy (anonymous):

Ellie’s bird feeder is shaped like a cone with a height of 24 in. and a radius of 3 in. Packages of bird seed are cylindrical and come in different sizes. What is the height of the package that has a radius of 6 in. and contains exactly enough bird seed to fill the feeder? Use 3.14 to approximate pi. in.

OpenStudy (anonymous):

@texaschic101 @Mehek14

Mehek (mehek14):

i gtg @texaschic101 will help

OpenStudy (anonymous):

@texaschic101 @amistre64 @Michele_Laino

OpenStudy (radar):

Volume of Bird Feeder=(1/3)Pi(3^2)24 volume of seed pkg= (pi)(6^2)h Simplify each Volume: 72pi (feeder) 36h(pi) seed pkg. The volumes need to be equal. 36 h (pi) = 72 Pi divide both sides by Pi 36 h = 72 solve for h

OpenStudy (anonymous):

2?

OpenStudy (radar):

Thats what I came up with.

OpenStudy (anonymous):

so 2 inches

OpenStudy (radar):

Yes the units would be inches. Note that the Pi(s) cancelled out in the division, so that would be an exact answer.